Assume treasuries are trading at 3% interest. ($100 investment that pays9 % interest after one year 6% interest after two years and 7 %interest after 3 years) what is the net present value?
Answer #1

The treasuries are trading at 3%. This would be the discount rate calculating the Present Value of Interest Income

For calculating the Present value, we will use the formula  

PV = Cash Flow / (1+ Discount Rate)^(Time), here time is in years from current time

Year Cash Flow Present Value of Cash Flow Details/ Comments
0 -$100 -$100 This is presumed to be the investment, which is a Cash outflow, and therefore a negative sign. A cash inflow would be considered with a positive sign
1 $9 $8.74

The interest earned on the investment = 100 (Investment) x 9% (Interest Rate) = $9

The PV of $9 earned after a year = 9/ (1+3%)^1 = $8.74

2 $6 $5.66

The interest earned on the investment in year 2 = 100 (Investment) x 6% (Interest Rate) = $6

The PV of $6 earned after 2 year = 6/ (1+3%)^2 = $5.66

3 $7 $6.41

The interest earned on the investment in year 3 = 100 (Investment) x 7% (Interest Rate) = $7

The PV of $7 earned after 3 year = 7/ (1+3%)^3 = $6.41

Net Present Value of Investment $79.20 NPV of Investment = Present Value of all Cash Flows (The Investment, a cash outflow and Earnings - Interest, the Cash Inflow)

The NPV of the Investment = $79.20